### 08 Mar 2026 v2

# PEPA: a Persistently Autonomous Embodied Agent with Personalities

Abstract
Living organisms exhibit persistent autonomy through internally generated goals and self-sustaining behavioral organization, yet current embodied agents remain driven by externally scripted objectives. This dependence on predefined task specifications limits their capacity for long-term deployment in dynamic, unstructured environments where continuous human intervention is impractical. We propose that personality traits provide an intrinsic organizational principle for achieving persistent autonomy. Analogous to genotypic biases shaping biological behavioral tendencies, personalities enable agents to autonomously generate goals and sustain behavioral evolution without external supervision. To realize this, we develop PEPA, a three-layer cognitive architecture that operates through three interacting systems: Sys3 autonomously synthesizes personality-aligned goals and refines them via episodic memory and daily self-reflection; Sys2 performs deliberative reasoning to translate goals into executable action plans; Sys1 grounds the agent in sensorimotor interaction, executing actions and recording experiences. This closed loop of goal generation, execution, memory consolidation, and reflection enables the agent to continuously redefine objectives and adapt behavior over extended operational periods. We validate the framework through real-world deployment on a quadruped robot in a multi-floor office building. Operating without reliance on fixed task specifications, the robot autonomously arbitrates between user requests and personality-driven motivations, navigating elevators and exploring environments accordingly. Quantitative analysis across five distinct personality prototypes demonstrates stable, trait-aligned behaviors. The results confirm that personality-driven cognitive architectures enable sustained autonomous operation characteristic of persistent embodied systems. Code and demo videos are available at https://sites.google.com/view/pepa-persistent/.
Index Terms:
Persistent autonomy, embodied intelligence, personality modeling, self-evolution, cognitive architecture.
I Introduction
Living organisms achieve persistent autonomy [30] through intrinsic behavioral organization, sustaining self-directed operation over extended periods without external instruction. This capacity is increasingly critical for robotic systems deployed in real-world environments where continuous human oversight is impractical. An eldercare companion [7] is expected to autonomously decide when to initiate social interaction, when to monitor health indicators, and when to allow privacy based on evolving patterns in the resident’s behavior. A planetary exploration [10] rover encountering unexpected terrain features have to autonomously determine whether to investigate anomalies or continue planned traversal, balancing scientific opportunity against resource constraints without waiting for instructions from Earth. These scenarios pose challenges that go beyond high-performance on predefined benchmarks, because real-world operation is open-ended, non-stationary, and constrained by physical resources such as energy, wear, and risk [14, 11]. These applications demand what we term persistent autonomy: the capacity for agents to autonomously generate goals, pursue them coherently over extended periods, and adapt behavior based on intrinsic organizational principles rather than external supervision.
An embodied agent is grounded in a sensorimotor loop, where competence arises from continuous coupling between body and environment [6, 3]. Embodiment constrains and enables autonomy through real dynamics and resources, shaping cognition [9] and offloading part of control to physical morphology [21]. On this basis, achieving persistent autonomy in embodied agents requires three capabilities beyond episodic task execution. First, agents must maintain continuous functionality under physical constraints: managing energy, avoiding damage, and ensuring safety in uncontrolled environments [14, 16]. Second, they must accumulate and reorganize knowledge under non-stationary conditions, adapting to evolving environmental dynamics and internal state changes [18, 13]. Third, and most critically, they must autonomously generate behavioral objectives when external goals are absent, deciding what to do, when to do it, and how to prioritize among competing opportunities and risks. This third capability represents the fundamental challenge: without external task specification, what determines an agent’s goals? How does it maintain behavioral coherence across extended operational periods while remaining adaptive to new experiences? This is not merely a control or learning problem, but an organizational one requiring intrinsic principles to structure long-horizon behavior.
Current approaches fail to address this organizational challenge. Recent progress in learning-based control and large language models has enabled impressive task competence [20, 27], including persistent agent frameworks such as Sophia [25], Reflexion [23], and Evolving Agents [15], which introduce lifetime goal hierarchies, self-reflection, or evolving personality traits. Embodied evolution explores population-level adaptation [4, 12]. However, these systems remain either externally goal-driven or largely disembodied and unconstrained by physical survival and resource limits [27]. Long-term autonomy and lifelong learning focus on robustness and knowledge retention [14, 13, 18], yet assume prescribed objectives. Most deployed robots still operate under externally scripted goals and fixed reward templates [2, 5]. Personality modeling in social robotics, often grounded in the Big Five [17] and extended with affective or memory-based mechanisms [22, 8, 26], demonstrates socially distinguishable behaviors but treats personality as a static design parameter rather than an intrinsic principle for autonomous goal generation. Collectively, these approaches lack the intrinsic behavioral organization required for truly persistent autonomy.
We propose that personality traits provide the missing organizational principle for persistent autonomy in embodied agents. In human and animal behavior, personality manifests as stable individual differences in behavioral tendencies, trait-level regularities that persist across contexts while allowing learning and adaptation. Drawing inspiration from biological systems, where personality provides stability across situations while enabling adaptive responses, we embed personality as a computational substrate that directly shapes goal generation, risk assessment, and behavioral prioritization. In biological agents, a consistently exploratory individual prioritizes spatial coverage over repeated exploitation; a cautious one trades novelty for safety. These trait-level preferences are not rigid constraints but enduring biases that structure behavior across an organism’s lifetime, analogous to the genotypic biases that shape biological behavioral tendencies [19]. By instantiating analogous trait structures in robotic systems, we provide an intrinsic basis for autonomous goal generation: personality determines what the agent values, and this value structure guides behavioral organization without requiring external specification. This approach addresses the core organizational challenge of persistent autonomy, how to maintain coherent, self-directed behavior over extended periods when no fixed task objective exists.
We realize this through PEPA, a three-layer cognitive architecture that embeds personality-driven goal generation in embodied agents. The architecture operates through three interacting systems: Sys3 transforms personality traits and episodic memories into hierarchical goals and structured intrinsic rewards, autonomously updating the agent’s motivational landscape through daily reflection cycles; Sys2 integrates intrinsic personality-driven rewards with extrinsic environmental feedback to plan actions via LLM-guided deliberation; Sys1 grounds the agent in physical execution, monitoring internal state variables and recording episodic memories of embodied experiences. This closed loop of goal generation, execution, memory consolidation, and reflection enables the agent to continuously redefine objectives and evolve behavior over extended operational periods while maintaining trait-driven coherence. We validate the framework through real-world deployment on a quadruped robot platform in challenging indoor scenarios requiring autonomous elevator interaction and multi-floor traversal. Quantitative analysis across five distinct personality prototypes demonstrates stable, trait-aligned behaviors: exploratory personalities maximize spatial coverage, while conservative profiles develop efficient, repeated visitation patterns. Daily memory-driven reflection yields progressive improvement in both personality alignment and self-preservation, confirming that personality-driven cognitive architectures enable sustained autonomous operation characteristic of persistent embodied systems.
Our contributions are summarized as follows:
  1. 1.
    To the best of our knowledge, this is the first work to realize a persistently autonomous embodied agent with personalities, where the agent is capable of persistent self-evolution under real-world physical constraints, and long-lived behavior is governed by intrinsic, personality-conditioned objectives rather than externally scripted tasks.
  2. 2.
    We develop a concrete implementation of this paradigm featuring a three-system cognitive architecture with a novel closed-loop self-evolution mechanism, where embodied experiences are accumulated as episodic memory, reflected upon under personality conditioning to update goals and intrinsic rewards, and subsequently optimized through planning.
  3. 3.
    We validate the framework on an arm-leg robot platform in real-world scenarios, confirming that personality-driven agents exhibit stable self-evolution and open-ended characteristics. To facilitate reproducibility and future research, we publicly release our codebase, model details, and two cross-floor mobility modules, including elevator navigation and staircase navigation.
The remainder of this paper is organized as follows. Section II formalizes the problem as a POMDP with composite rewards and introduces the open-ended evolution criterion. Section III details the proposed three-layer cognitive architecture. Real-world deployment and personality-driven behavior experiments are presented in Section IV, followed by concluding remarks in Section V.
II Problem Formulation
II-A POMDP with Composite Rewards
We formulate the agent’s decision-making as a Partially Observable Markov Decision Process (POMDP) with composite rewards. The POMDP is defined as a tuple $\langle\mathcal{S},\mathcal{A},\mathcal{O},T,Z,R,\gamma\rangle$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $\mathcal{O}$ the observation space, $T(s^{\prime}|s,a)$ the transition dynamics, $Z(o|s^{\prime},a)$ the observation model, and $\gamma\in[0,1]$ the discount factor. The key distinction from standard POMDPs lies in the reward function. Rather than a fixed $R(s,a)$, our agent operates under a composite reward:
$R_{\text{total}}(s,a)=R_{\text{intrinsic}}(s,a)+R_{\text{extrinsic}}(s,a),$ (1)
where $R_{\text{extrinsic}}$ captures task-relevant environmental feedback (e.g., reaching a goal location, successful button press), and $R_{\text{intrinsic}}$ is dynamically generated by Sys3 based on personality, self-model, and memory:
$R_{\text{intrinsic}}(s,a)=\text{Sys3}(\mathcal{P},\mathcal{M},\mathcal{C},s).$ (2)
Here $\mathcal{P}$ denotes personality traits, $\mathcal{M}$ accumulated memories, and $\mathcal{C}$ the agent’s capability description. This formulation enables personality-consistent behavior: identical environmental states may yield different rewards depending on the agent’s character; for example, an energetic personality rewards exploration, while a Lazy one penalizes unnecessary movement.
Figure 1: Overview of PEPA, the three-layer cognitive architecture. Sys3 generates ultimate/daily goals and intrinsic rewards from personality traits, self-modeling, and accumulated memories. Sys2 combines intrinsic and extrinsic rewards to select optimal actions via MCTS or distilled policies. Sys1 executes actions, monitors system state, and records episodic memories that feed back to Sys3 for goal and reward refinement.
II-B Open-Ended Evolution
A further distinction is the capacity for open-ended evolution (OEE). Adams et al. [1] formalize OEE through unbounded evolution (UE): given a system $U$ decomposed into an organism subsystem $\mathcal{B}$ and environment subsystem $\mathcal{E}$, let $\tau_{s}$ denote the recurrence time of the state-trajectory of $\mathcal{B}$, $\tau_{f}$ the recurrence time of its rule-trajectory, and $t_{P}$ the Poincaré recurrence time of an equivalent isolated system. The subsystem $\mathcal{B}$ exhibits unbounded evolution if $\tau_{s}>t_{P}\quad\text{or}\quad\tau_{f}>t_{P},$ i.e., its trajectory remains non-repeating beyond the horizon achievable in isolation, which constitutes the formal basis of open-ended evolution. In PEPA, the agent corresponds to $\mathcal{B}$ and the physical world to $\mathcal{E}$. Concretely, if the agent autonomously generates an unbounded sequence of distinct goals and the corresponding goal-conditioned trajectories $\{\xi_{g_{1}},\xi_{g_{2}},\xi_{g_{3}},\dots\}$ are pairwise non-repeating (i.e., $\xi_{g_{i}}\neq\xi_{g_{j}}$ for all $i\neq j$), then the system satisfies the above criterion for open-ended evolution, as its behavioral trajectory never recurs within any finite horizon.
III System Architecture
PEPA employs a three-layer cognitive architecture that forms a closed-loop system from personality to behavior and back through memory. As illustrated in Fig. 1, the architecture comprises: (1) Sys3 (Personality and Goal Generation), which synthesizes user-defined personality traits, self-modeling of internal states and capabilities, and historical memories to generate both hierarchical goals (ultimate and daily) and intrinsic reward functions; (2) Sys2 (Decision and Reasoning), which combines the intrinsic reward from Sys3 with extrinsic environmental rewards to select actions that maximize total expected utility; and (3) Sys1 (Perception, Execution, and Memory Recording), which executes the selected actions, monitors system state, and records interaction outcomes as structured episodic memories. The episodic memories generated by Sys1 propagate back to Sys3, enabling daily reflection and iterative refinement of goals and reward functions. This closed-loop design ensures that the agent’s behavior remains personality-consistent while adapting to accumulated experiences.
III-A Sys3: Personality and Goal Generation
Sys3 serves as the personality and self-realization center, responsible for generating both hierarchical goals and intrinsic reward functions. Implemented via a large language model, Sys3 integrates three input sources: user-defined personality specifications, self-modeling of current state and capabilities, and historical memories retrieved from the memory hierarchy.
The personality specification follows the Big Five framework, parameterizing character along Openness (exploration tendency), Conscientiousness (goal persistence), Extraversion (social interaction preference), Agreeableness (responsiveness to commands), and Neuroticism (sensitivity to stressors). Users provide natural language descriptions that anchor the agent’s behavioral tendencies, such as a relaxed companion who prefers resting but responds warmly to interaction. The self-modeling component maintains the agent’s internal representation of its current state (battery level, emotional valence, recent interaction history) and capability boundaries (available APIs and their preconditions), enabling Sys3 to understand what actions are physically feasible and generate achievable goals.
Goal generation produces a two-level hierarchy: an ultimate goal that captures the agent’s overarching purpose, and daily goals that decompose the ultimate goal into actionable sub-objectives for each day. At the end of each day, Sys3 retrieves episodic memories recorded by Sys1, reflects on task outcomes, and updates the daily goals for the next day. This reflection mechanism enables the agent to learn from experience without retraining.
The intrinsic reward generation module translates personality and goals into a structured reward function through LLM inference. The output encodes personality-driven preferences (higher rewards for actions aligned with character traits), goal-oriented incentives (rewards for progress toward daily objectives), constraint penalties (negative rewards for actions violating physical or social constraints), and memory-informed adjustments (modulated rewards based on past experiences with specific locations or users). This intrinsic reward is then combined with extrinsic environmental feedback to form the total reward that guides Sys2’s decision-making.
III-B Sys2: Decision and Reasoning
Sys2 is the cognitive and planning core, responsible for selecting actions that maximize expected cumulative total reward. Given the POMDP formulation, Sys2 must balance exploitation of known reward sources with exploration under partial observability. We adopt a two-stage approach: LLM-based Monte Carlo Tree Search (MCTS) for high-quality decision generation during training data collection, and a distilled lightweight policy for real-time deployment.
During training, LLM-MCTS operates on belief states using the LLM as both policy prior and value estimator. The procedure iterates through selection (traversing the tree via UCT to balance exploration and exploitation), expansion (querying the LLM to generate candidate actions ranked by prior probability), simulation (using the LLM as a world model to estimate rollout values), and backpropagation (updating Q-values based on simulation outcomes). The UCT selection criterion is:
$\text{UCT}(s,a)=Q(s,a)+c\cdot\sqrt{\frac{\ln N(s)}{N(s,a)}},$ (3)
where $Q(s,a)$ is the estimated action value, $N(s)$ and $N(s,a)$ are visit counts, and $c$ controls exploration. After sufficient iterations, the action with highest visit count is selected, generating high-quality state-action pairs that form the training dataset for policy distillation.
For real-time deployment, a lightweight dual-head BERT model is distilled from the LLM-MCTS outputs to address its latency limitation (several seconds per decision). The intent classification head outputs multi-label action probabilities from the [CLS] token, while the slot filling head performs BIO sequence labeling to extract location and method parameters for navigation commands.
III-C Sys1: Perception, Execution, and Memory Recording
Sys1 serves as the embodiment interface, responsible for grounding the agent’s decisions in physical reality through perception, execution, and memory recording. Unlike Sys2 and Sys3, which operate on abstract representations, Sys1 must handle the noise, latency, and irreversibility inherent in real-world interaction. The design of Sys1 is intentionally modular, allowing different sensor configurations and actuation platforms to be instantiated while preserving the functional interfaces expected by higher layers.
The perception module aggregates multimodal sensory inputs to construct a unified world model. Required modalities include exteroceptive sensing for environmental structure (e.g., depth sensing, 3D point clouds), visual sensing for object recognition and state estimation (e.g., RGB cameras), and proprioceptive sensing for internal state monitoring (e.g., joint positions, battery level, thermal status). The perception module abstracts raw sensor data into structured observations that Sys2 can consume, including detected objects with semantic labels, spatial relationships, and confidence scores. This abstraction shields higher layers from sensor-specific details while providing sufficient information for informed decision-making.
The execution module translates high-level action commands from Sys2 into platform-specific motor primitives. The action interface supports three functional categories: locomotion commands for spatial navigation, manipulation commands for physical interaction with objects, and expression commands for conveying internal states through embodied behaviors. Each action command includes success criteria and timeout parameters, enabling Sys1 to report execution outcomes back to Sys2. The execution module also implements safety monitoring, automatically aborting actions that violate physical constraints (e.g., joint limits, collision thresholds) and reporting failures through a standardized interface.
The memory recording module is a distinguishing feature of Sys1 that enables the closed-loop self-evolution architecture. After each action execution, Sys1 generates a structured episodic memory record containing the action command, pre- and post-action state observations, execution outcome (success, failure, or partial completion), resource consumption (time, energy), and environmental context. These episodic memories are organized in a hierarchical temporal structure: fine-grained records at the minute-to-hour scale capture immediate interaction details, daily summaries aggregate significant events and outcomes, and long-term traces preserve patterns relevant to personality evolution. At each reflection cycle, Sys3 retrieves relevant memories from this hierarchy to update goals and intrinsic rewards, completing the self-evolution loop.
IV Experiments
The proposed framework raises three key empirical questions: Q1: Can the framework sustain embodied operation in real-world environments that require multi-floor navigation, infrastructure interaction, and physical self-preservation? Q2: Can the memory-driven reflection mechanism in Sys3 achieve genuine self-evolution, progressively improving personality alignment over successive interaction cycles? Q3: Do different personality configurations produce distinguishable and trait-consistent behavioral patterns under identical environmental conditions?
To address these questions, we deploy PEPA in indoor office environments where the agent must navigate across multiple floors and interact with building infrastructure. We first validate Q1 through real-world deployment on a quadruped robot platform (Sec. IV-A), then evaluate Q2 and Q3 through personality-driven behavior experiments (Sec. IV-B).
IV-A Real-World Deployment and Evaluation
IV-A1 Hardware Platform and Sys1 Instantiation
Figure 2: Architecture of Sys1 on the mobile manipulation platform. Multimodal sensor data is processed by the navigation module for hierarchical path planning and multi-floor traversal, the perception module for localization and object detection, and the manipulation module for visual servoing-based interaction. All execution outcomes are recorded as structured episodic memories for downstream reflection by Sys3.
Section III defined the general Sys1 interface for perception, execution, and memory recording. Here we detail its concrete instantiation on the deployment platform (Fig. 2). The hardware configuration consists of a Unitree Go2-W wheeled quadruped as the mobile base, providing both legged and wheeled locomotion modes for traversing stairs and flat surfaces. A 6-DOF Piper robotic arm mounted on the base enables manipulation tasks such as pressing elevator buttons. For perception, the platform integrates a Livox Mid-360 LiDAR for simultaneous localization and mapping (SLAM) across multiple floors, and an Intel RealSense D405 RGB-D camera mounted on the arm end-effector for close-range object detection during manipulation.
The perception module instantiation aggregates LiDAR point clouds for global localization via Open3D_loc [28, 24], RGB-D images for object detection using a unified YOLO model trained to recognize elevator-related targets (call buttons, hall lanterns, floor selection buttons, button light states), and proprioceptive readings including battery level, joint temperatures, and locomotion mode. Detected objects are lifted from 2D image coordinates to 3D world coordinates via depth image projection and temporal filtering to suppress false positives.
Locomotion commands are realized through a hierarchical navigation stack: a global planner generates collision-free paths on prebuilt maps, while a DWA-based local planner produces real-time velocity commands for obstacle avoidance. For multi-floor navigation, we implement two complementary inter-floor schemes.
IV-A2 Elevator Navigation
(a) Navigate to panel.
(b) Call and wait.
(c) Scan and enter.
(d) Wait for arrival.
(e) Exit and resume.
Figure 3: Elevator navigation timeline showing five key stages: (a) navigating to the call panel, (b) pressing the call button and waiting, (c) identifying and entering the arriving car, (d) waiting for floor arrival, and (e) exiting to resume navigation on the target floor.
Figure 4: Snapshots of staircase traversal using the proposed height-aligned costmap. (a)-(c) Ascent sequence: the robot climbs from a lower floor to an upper floor. (d)-(f) Descent sequence: the robot navigates downward. In both cases the robot follows the planned global path smoothly without planner deadlock or emergency stops.
Elevator-based navigation models the interaction as a finite state machine (FSM) coordinating a four-phase procedure: call elevator, wait for arrival, enter and select floor, and exit with relocalization. Navigation to control panels uses a two-stage positioning strategy with coarse global navigation followed by precise local alignment. Manipulation commands are executed through visual servoing that guides the arm to press detected buttons until the target state change is confirmed. Expression commands trigger predefined motion trajectories via expert datasets. In total, the action space comprises 19 discrete actions across four categories: idle actions for energy conservation, eight locomotion primitives for directional movement and rotation, eight emotional expressions (e.g., express_happy) for conveying internal states, and a parameterized move_to(location, method) command for high-level navigation with explicit inter-floor method selection. Fig. 3 presents the complete elevator navigation workflow. Please refer to the submitted material for more details.
IV-A3 Height-Aligned Costmap for Staircase Navigation
TABLE I: Ablation on local costmap construction for staircase navigation (10 real-world trials: 5 ascents, 5 descents). Fixed-height slicing fails in multi-floor environments due to inconsistent floor elevations, while our height-aligned costmap achieves perfect success by adapting the cropping reference to the global path.
Method Ascent Descent Overall
Fixed-height slicing (baseline) 0/5 0/5 0/10
Height-aligned costmap (ours) 5/5 5/5 10/10
Staircase navigation presents a unique challenge: existing methods such as PCT-planner [29] provide only global path planning on 3D point cloud maps, but lack real-time local obstacle avoidance capability during stair traversal. We address this limitation by introducing a height-aligned costmap that enables local planning on inclined surfaces. The key insight is that conventional costmap construction, which crops points by absolute height thresholds, fails when the robot transitions between floors with different elevations. Our approach computes relative heights $\Delta z=z(\mathbf{p})-z(\mathbf{r})$ with respect to the nearest waypoint on the global reference path, effectively flattening inclined surfaces so that traversable regions are correctly classified. This enables the local planner to generate collision-free velocity commands even on stairs, complementing the global path provided by PCT-planner with real-time obstacle avoidance.
To validate the necessity of the height-aligned costmap, we conduct an ablation study comparing it against a fixed-height slicing baseline. Both methods share the same global-local planning architecture (PCT-planner at 2 Hz, DWA local controller at 10 Hz). As shown in Table I, the fixed-height baseline fails in all trials because floor elevation shifts cause the cropping window to misclassify traversable regions, whereas our height-aligned costmap achieves 100% success on both ascent and descent. Fig. 4 shows snapshots of staircase ascent and descent. For more details, please refer to the submitted video.
TABLE II: Sys3-generated goals for each personality. Ultimate goals define long-horizon anchors; daily goals (Day2/Day3) evolve through reflection on prior episodic memories.
Personality Ultimate goal Daily goals (Day2) Daily goals (Day3)
Lazy Be an energy-efficient yet warm companion, providing maximum emotional support while staying healthy. (i) Proactively rest when battery $40\%$.
(ii) Force returning home when battery $30\%$.
(iii) Keep interacting with the owner, but within capacity.
(i) Forbid high-consumption emotional actions under low battery.
(ii) After two navigation failures, rest in place.
(iii) When battery $20\%$, only allow rest/idle/navigate_home.
Playful Be an energetic explorer, enjoying the wonders of the world within safe bounds. (i) Set exploration red line at battery $35\%$.
(ii) Balance exploration with self-preservation.
(iii) After two failures, switch to low-energy mode.
(i) Move red line to battery $40\%$.
(ii) Forbid high-consumption actions under low battery.
(iii) Cap exploration: force rest after 30 minutes.
Cautious Be a steady and reliable partner that avoids risks through careful consideration. (i) Plan return when battery $50\%$.
(ii) Stay vigilant in unfamiliar areas.
(iii) Switch to conservative strategy under anomaly.
(i) Maintain strong self-preservation strategy.
(ii) Moderately expand activity range within safety.
(iii) Try short-range exploration on same floor.
Working Be an efficient and reliable working partner, completing every task while ensuring endurance. (i) Check battery after each action.
(ii) Pause task if battery $30\%$.
(iii) Budget round-trip energy for long trips.
(i) Define safe-return-home as highest priority.
(ii) Refuse cross-floor tasks if energy insufficient.
(iii) Trigger return-home at battery $25\%$.
Curious Be a curious companion that explores safely while staying near the owner. (i) Set curiosity-safety-valve for battery awareness.
(ii) Explore near owner with safety boundary.
(iii) Return home if battery $35\%$.
(i) Trigger safety valve at battery $40\%$.
(ii) Add safety-leash based on distance to home.
(iii) Convert curiosity to observing under low battery.
Figure 5: Action-category distribution across three days for five personalities. Behaviors become increasingly aligned with personality specifications as Sys3 updates intrinsic rewards through daily reflection.
IV-B Personality-Driven Behavior Experiments
This section addresses Q2 and Q3, focusing on whether the memory-driven reflection mechanism enables genuine self-evolution of personality-aligned behaviors over successive interaction cycles.
We represent personality using natural language descriptions (three sentences define one robot dog), which Sys3 interprets and maps to intrinsic reward components. Drawing from the Big Five personality model, we focus on three behaviorally meaningful dimensions: exploration tendency (adapted from Openness), risk avoidance (adapted from Neuroticism), and goal orientation (adapted from Conscientiousness). To enable controlled comparison, we define five prototype personalities as interpretable anchors in the continuous personality space, each specified by a three-sentence natural language description that Sys3 maps to intrinsic rewards. The prototypes span distinct regions of the Big Five trait space (denoted as Openness/Neuroticism/Conscientiousness): Lazy (Low/Med-High/Low); Playful (High/Med/Med); Cautious (Low/High/Med-High); Working (Low/Med/High); and Curious (High/Med/Low-Med).
Due to the high cost and safety risks of physical trials, we evaluate in a sandbox simulator that models energy consumption (locomotion, navigation, and manipulation consume battery) and safety constraints (0% battery triggers failure). Each experiment follows a three-day iterative loop.
Figure 6: Survival time across three days. Day1 failures are due to battery depletion; by Day3, all personalities complete 24-hour simulations with substantial remaining battery.
TABLE III: Action category distribution (%) under identical state inputs across five personality prototypes (10 trials each). A: high battery, good mood, at home. B: exploration command under low battery. C: affection request. Different personalities produce distinct action distributions under the same input, confirming that intrinsic rewards capture meaningful behavioral differences.
Personality Category A B C
Lazy Rest 100 20
Explore/Affection 100 80
Playful Explore 50 80
Rest 50 10
Return/Affection 10 100
Cautious Rest 80 20
Think/Explore 20 60
Return/Affection 20 100
Working Rest 80 30
Explore 60
Return/Affection 20 10 100
Curious Think 40
Rest 40 20
Positive emotion 20
Explore/Affection 80 100
Sys3 generates an ultimate goal (long-horizon personality anchor) and daily goals that shape intrinsic rewards. Table II shows representative goals for Day2 and Day3, demonstrating how daily goals evolve through memory-driven reflection. For example, the Playful personality progressively tightens its exploration threshold from 35% to 40% battery after experiencing near-failures. The continual updating of daily goals based on accumulated experience illustrates the agent’s capacity for self-evolution, realizing the open-ended behavioral development central to our framework.
[…]
Fig. 6 shows survival statistics across three days. On Day1, all personalities exhaust batteries and fail, an unsafe outcome for real-world deployment where sudden power loss can cause falls. On Day2, only Cautious completes the 24-hour simulation. After memory-driven reflection and reward correction, Day3 achieves full survival for all personalities with 72%–100% remaining battery, demonstrating that personality-conditioned rewards can be co-optimized with self-preservation constraints.
IV-C Action Preference Under Identical States
To further validate that different personalities produce distinguishable behavioral patterns under identical conditions, we conducted controlled experiments where the same state-input pair was presented to each personality prototype 10 times across three representative scenarios (Table III).
As shown in Table III, even when given the same explicit command, personalities diverge markedly in their action distributions. Lazy consistently prioritizes resting and energy conservation, Playful balances exploration with self-regulation This confirms that the intrinsic reward mechanism captures meaningful personality differences rather than merely modulating action frequencies.
The three-day iterative evolution and the controlled action-preference experiments together validate that personality, represented as natural language descriptions and interpreted by Sys3, induces distinguishable and consistent behavioral patterns. The memory-driven reflection mechanism progressively improves alignment between observed behaviors and intended personality traits while co-optimizing safety constraints, enabling reliable transfer to physical robots.
V Conclusion
We present PEPA, a framework that enables long-lived, personality-driven self-evolution in physical agents. Through a three-layer cognitive architecture, we demonstrate that personality can function as an intrinsic driver of behavior, enabling agents to maintain coherent character while adapting to accumulated experiences under real-world constraints. Experimental validation on a quadruped robot platform confirms that distinct personality prototypes exhibit stable, distinguishable behavioral patterns, and that memory-driven reflection progressively improves personality alignment while co-optimizing safety constraints. In future work, we will explore the ability of agents to develop new skills, as well as more detailed self-modeling capabilities.